### v1

# Steel-LLM: From Scratch to Open Source – A Personal Journey in Building a Chinese-Centric LLM

Abstract
Steel-LLM is a Chinese-centric language model developed from scratch with the goal of creating a high-quality open-source model despite limited computational resources. Launched in March 2024, the project aimed to train a 1-billion-parameter model on a large-scale dataset, prioritizing transparency and the sharing of practical insights to assist others in the community. The training process primarily focused on Chinese data, with a small proportion of English data included, addressing gaps in existing open-source LLMs by providing a more detailed and practical account of the model-building journey. Steel-LLM has demonstrated competitive performance on benchmarks such as CEVAL and CMMLU, outperforming early models from larger institutions. This paper provides a comprehensive summary of the project’s key contributions, including data collection, model design, training methodologies, and the challenges encountered along the way, offering a valuable resource for researchers and practitioners looking to develop their own LLMs. The model checkpoints and the training script are available at https://github.com/zhanshijinwat/Steel-LLM.
Qingshui Gu
Tsinghua University
gqs19@tsinghua.org.cn
Tianyu Zheng
Beijing University of Posts
and Telecommunications
zhengtianyu@bupt.edu.cn
Shu Li
Tsinghua University
lishu14@tsinghua.org.cn
Zhaoxiang Zhang
Institute of Automation,
Chinese Academy of Sciences
zhaoxiang.zhang@ia.ac.cn
1 Introduction
The rapid advancements in open-source large language models (LLMs) have led to significant achievements in natural language processing (NLP), enabling applications ranging from conversational agents to code generation. However, despite these advances, challenges remain in the transparency, accessibility, and resource efficiency of LLM development. Many prominent LLMs, such as the Qwen series (Bai et al., 2023a; Yang et al., 2024), Llama (Touvron et al., 2023b), and Deepseek (DeepSeek-AI et al., 2024a), provide only the final model weights while withholding essential details such as training data, code, and intermediate checkpoints. This limited openness creates obstacles to reproducibility and prevents the broader research community from building upon these models effectively.
Several initiatives, such as LLM360’s Amber (Liu et al., 2023; Tan et al., 2024), M-A-P’s Neo (Zhang et al., 2024b), and AI2’s OLMO series (Groeneveld et al., 2024a; OLMo et al., 2024; Muennighoff et al., 2024), have addressed these limitations by releasing complete training pipelines, datasets, and intermediate checkpoints. While these contributions are invaluable, their development typically demands extensive computational resources, making them inaccessible to smaller research teams or individual practitioners. This creates a significant gap for fully transparent and resource-efficient LLMs tailored for smaller-scale research efforts, particularly in non-English languages such as Chinese.
This paper introduces Steel-LLM, a small, fully open-source, Chinese-centric LLM developed with limited computational resources. Launched in March 2024, the project aimed to train a 1-billion-parameter model on a large-scale dataset, prioritizing transparency and the sharing of practical insights to assist others in the community. Unlike many large-scale projects, Steel-LLM demonstrates that high-quality LLMs can be developed efficiently while maintaining transparency. This work focuses on the Chinese language, with a small proportion of English data, addressing gaps in existing open-source LLMs by providing a more detailed and practical account of the model-building journey.
Steel-LLM achieves competitive performance on benchmarks such as CEVAL and CMMLU, outperforming early models from larger institutions. The model architecture incorporates innovative features such as Soft Mixture of Experts (Soft MOE) and an enhanced Feed-Forward Network, optimizing performance within resource constraints. Our training framework, modified from TinyLlama (Zhang et al., 2024c), includes several optimizations for efficiency and usability, such as improved model loading, training progress restoration, and the ability to append data during the training process.
The contributions of this work are as follows:
  • Resource-Efficient Model Development: We present a 1-billion-parameter model trained primarily on Chinese data, with a small proportion of English, addressing the need for more diverse language representation in open-source LLMs. By leveraging limited computational resources (8 GPUs), we demonstrate that high-quality LLMs can be developed without access to large-scale infrastructure.
  • Complete Transparency: We offer complete transparency in our development process, including the release of our training pipeline, dataset, model architecture, and intermediate checkpoints. This facilitates reproducibility and enables further research by the broader community.
  • Practical Guidance for Small-Scale Research: We provide detailed insights into our model architecture, training framework, and data preparation process, offering practical guidance for researchers and practitioners with limited resources. This includes optimizations for training efficiency, such as mixed-precision training, FlashAttention, and operator fusion.
  • Benchmark Performance: Steel-LLM achieves competitive performance on Chinese benchmarks such as CEVAL and CMMLU, outperforming some early models from larger institutions. This demonstrates the effectiveness of our approach in developing a high-quality Chinese-centric LLM with limited resources.
By prioritizing resource efficiency, openness, and practical applicability, this work contributes to the broader LLM research community, offering a valuable resource for replicating or extending similar efforts with fewer computational constraints. All model checkpoints, training scripts, and related resources are fully open-sourced, further promoting transparency and collaboration in the field of LLM development.
2 Related Works
Recent developments in open-source LLMs have varied widely in terms of transparency and accessibility. Many models, such as Qwen (Bai et al., 2023a; Yang et al., 2024), Llama (Touvron et al., 2023b), Deepseek (DeepSeek-AI et al., 2024a), Gemma  (Team et al., 2024), InternLM  (Cai et al., 2024), Mixtral  (Jiang et al., 2023), Yi  (AI et al., 2024), GLM  (GLM et al., 2024), have been released with only the final model checkpoints and weights, while withholding crucial details such as training data, codes, and intermediate checkpoints. This limited transparency hinders reproducibility and makes it difficult for the broader research community to fully understand or build upon these models.
In response to these challenges, several initiatives have adopted a more open approach by releasing complete training pipelines, datasets, and intermediate checkpoints. Notable examples include LLM360’s Amber (Liu et al., 2023; Tan et al., 2024), M-A-P’s MAP-Neo (Zhang et al., 2024b), and AI2’s OLMO series (Groeneveld et al., 2024a; OLMo et al., 2024; Muennighoff et al., 2024), all of which offer comprehensive resources, including training code, model weights, and intermediate checkpoints. While these contributions are invaluable to the field, the large-scale nature of these models necessitates significant computational resources, which may render them inaccessible to smaller research teams or individual practitioners. Consequently, there remains a gap for open-source LLMs that are both fully transparent and feasible for smaller teams to develop and deploy.
This paper addresses this need by presenting a small, fully open-source, Chinese-centric LLM. Designed to minimize resource requirements, our model offers a complete end-to-end solution for researchers with limited computational access. By sharing the full training pipeline, dataset, and model architecture, we aim to make the development of high-quality LLMs more accessible. This work emphasizes transparency and practical guidance, providing a valuable resource for those seeking to replicate or extend the model with fewer computational constraints.
3 Architecture
The model structure of Steel-LLM is adapted from Qwen(Bai et al. (2023a)).A Transformer block can be roughly divided into self-attention and Feed-Forward Network (FFN). An efficient implementation of self-attention is Flash Attention(Dao et al. (2022)), which is widely utilized. Flash Attention not only improves the efficiency of model training and inference but also saves GPU memory. Steel-LLM reuses Flash Attention and only makes improvements to the FFN layer. In the FFN layer, we adopts Soft Mixture of Experts (Soft MOE, Puigcerver et al. (2024)) and enhances the second layer of the MLP. The architecture of Steel-LLM’s Transformer block is illustrated in Figure 1.
Figure 1: The architecture of Steel-LLM Transformer block
3.1 Soft MOE
Mixture of Experts (MOE) (Jacobs et al., 1991) was first proposed in 1991 and is widely used in the field of recommendation systems  (Ma et al., 2018). In the architecture of large language models, sparse MoE is commonly employed, such as in deepSeekMoE(Dai et al. (2024)), Qwen-MoE(Team (2024c)), etc. A typical practice to construct an MoE language model usually replaces FFNs in a Transformer with MoE layers(Lepikhin et al. (2021);Zoph (2022);Fedus et al. (2022)). An MoE layer consists of multiple experts, and each expert is a FFN. A gating network calculates scores to assign each token to a small subset of experts. Under the condition of the same parameter scale, the sparse MOE model activates fewer parameters than a dense model, thus reducing the number of FLOPs.
Steel-LLM is trained using only on 8 GPUs with limited GPU memory. For a model featuring sparse MoE structure, all parameters must be loaded into the GPU memory. Owing to the sparse nature of expert selection, in contrast to a dense model, the FFN parameters of the sparse MoE model cannot be fully trained. Steel-LLM is a small-scale language model with 1 billion parameters, leading to a relatively low computational load. Meanwhile, our objective is to fully train each parameter and leverage the advantages of the MoE structure to enhance model performance. Consequently, we ultimately opt for the soft MoE(Puigcerver et al. (2024)) structure. Soft MoE is fully differentiable. Therefore, there is no need to consider problems such as expert unbalance that exist in sparse MoE.
We denote the input tokens for a single sequence as $\mathbf{X}\in\mathbb{R}^{m\times d}$, where $m$ represents the number of tokens and $d$ denotes their dimensions. In the soft MoE layer, each expert processes $p$ slots, and each slot has a corresponding d-dimensional learnable parameter matrix $\Phi\in\mathbb{R}^{d\times(n\cdot p)}$. The number of slots is a crucial hyperparameter for adjusting the time complexity of the soft MoE layer. the input slots $\tilde{\mathbf{X}}\in\mathbb{R}^{(n\cdot p)\times d}$ are obtained through convex combinations of all the $m$ input tokens, which can be computed as follows:
$\mathbf{D}_{ij}=\frac{\exp((\mathbf{X}\bm{\Phi})_{ij})}{\sum_{i^{\prime}=1}^{m}\exp((\mathbf{X}\bm{\Phi})_{i^{\prime}j})},\quad\tilde{\mathbf{X}}=\mathbf{D}^{\top}\mathbf{X}$
We denote $D$ as the dispatch weight matrix. It is obtained by applying the softmax function column-wise to the matrix $\mathbf{X}\mathbf{\Phi}$. Then, the corresponding expert function is applied to each slot (i.e., on rows of $\tilde{\mathbf{X}}$) to obtain the output slots $\tilde{\mathbf{Y}}_{i}=f_{\lfloor i/p\rfloor}(\tilde{\mathbf{X}}_{i})$. The combination process is then carried out as follows:
$\mathbf{C}_{ij}=\frac{\exp((\mathbf{X}\bm{\Phi})_{ij})}{\sum_{j^{\prime}=1}^{n\cdot p}\exp((\mathbf{X}\bm{\Phi})_{ij^{\prime}})},\quad\mathrm{~{}}\mathbf{Y}=\mathbf{C}\tilde{\mathbf{Y}}$
We refer to $C$ as the combine weights, which is the result of applying the softmax function row-wise to the matrix $\mathbf{X}\mathbf{\Phi}$. The output tokens $Y$ are computed as a convex combination of all $(n\cdot p)$ output slots.
3.2 Enhanced Feed-Forward Network
Since the Steel-LLM employs the soft MoE approach, within this framework, the Feed-Forward Network (FFN) effectively represents an expert. In a vanilla Transformer, The FFN comprises two layers of Multi-Layer Perceptrons (MLPs). In the architecture of large language models, a prevalent strategy is to enhance the first layer of the MLP using the SwiGLU(Shazeer (2020)) activation function(Bai et al. (2023a);Touvron et al. (2023a);DeepSeek-AI et al. (2024b)). The SwiGLU activation function enhances the model’s non-linear representational capabilities, thereby improving its performance. Additionally, we extended the application of the SwiGLU activation function to the second layer of the MLP within the FFN.
Regarding other architectural elements, Steel-LLM adopts the modifications made by Qwen(Bai et al. (2023a)) to the transformer block. These modifications are widely used in open-source models such as LLama, Mixtral, and Deepseek.
Parameters Value
Layers 18
Heads 32
KV heads 32
Num_experts 6
Slots_per_expert 1
Hidden size 1,792
Intermediate size 1,792
Vocab size 151,936
Table 1: Key model parameters.
• Positional embedding.We intend to use the Rotary Position Embedding (RoPE)(Su et al. (2023)) for Steel-LLM. RoPE is a relative position encoding technique. Its core idea is to encode absolute position information into a rotation matrix, thereby representing the relative position relationships among tokens. During the training process, we adopt a global training precision of BF16, while for RoPE, we employ local FP32 precision.
• Bias. Most layers of Steel-LLM have no bias, except for the QKV layer(Chowdhery et al. (2022)).
• Pre-Norm & RMSNorm. Pre-normalization improves training stability compared to post-normalization. We normalize the inputs of self-attention layers and FFN layers with RMSNorm(Zhang & Sennrich (2019)).
The hyperparameters of Steel-LLM are presented in Table 1. We employed the Qwen1.5 tokenizer (Team (2024b)), which utilizes byte pair encoding (BPE).
4 Training Framework
Steel-LLM is trained using only 8 NVIDIA A100/H800 GPUs, so it is essential to maximize the training speed. Moreover, considering the potential reuse of our training code by others, we have made some usability optimizations.
Our pretraining framework, which is modified from TinyLlama(Zhang et al. (2024c)), has undergone the following optimizations.
• Model loading. In the open-source community, the file format of LLM usually follows the standards of the Transformers(Wolf et al. (2020)) library and it can be easily loaded through the AutoModelForCausalLM.from_pretrained function. To enhance usability and facilitate model loading for different architectures, our framework supports the Transformer library’s model loading method. The original framework, by contrast, was designed for training standard Llama architecture.
• Training Progress Restoration. In the training process, we are required to save not only the checkpoints of both the model and the optimizer but also the training progress of the data. The original framework’s method of saving the training step is the simplest for recording data progress, provided that there is no data change (no addition or deletion). This limitation restricts the flexibility of data organization during the training process. We choose to serialize the entire data management class PackedDatasetIterator using the pickle library and then save it. This includes the file names of all training data, the index of each data piece, etc.
• Appending Data During the Training Process. Training LLMs typically spans dozens of days. During this training, appending new data is a common practice. The simplest implementation approach is to train the newly added data at the end. Nevertheless, to safeguard against the significant distributional disparities between new and old data that could impact the model’s training performance, we have devised a method for re-shuffling the indices of both the newly appended data and the hitherto untrained data from the old dataset. Additionally, to avert the inadvertent repeated addition of data files to the training process, Steel-LLM has incorporated a function that utilizes a hash value, specifically the MD5(Rivest (1992)) hash algorithm, to detect data content duplication.
To imporve training speed and conserve GPU memory, during the training process, we utilized techniques such as bfloat16 mixed-precision(Kalamkar et al. (2019)), Fully Sharded Data Parallel (FSDP, Zhao et al. (2023)), and FlashAttention(Dao et al. (2022)). Operator fusion represents an additional training optimization approach. By integrating multiple computational steps, it reduces intermediate activations and memory access, and can be implemented via CUDA or Triton. Specifically, during training, we adopted the CUDA-based version of Rotary Position Embedding (RoPE,Su et al. (2023)) and the Triton-based(Tillet et al. (2019)) cross entropy loss function. We then conducted an ablation study using a micro-batch size of 8 and a 1.8-billion-parameter model on NVIDIA A100 GPU to analyze the impact of these techniques on training efficiency and GPU memory usage, as presented in Table 2. Employing all training acceleration techniques, the training speed can be enhanced by approximately 50%.
Exp 1 Exp 2 Exp 3 Exp 4 Exp 5 Exp 6 Exp 7 Exp 8
FlashAttention $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
SelfAttention(PyTorch) $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
RoPE(CUDA) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
RoPE(PyTorch) $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
RMSNorm(CUDA) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
RMSNorm(PyTorch) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$
Loss Function(Triton) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
Loss Function(PyTorch) $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
FSDP $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
FSDP(no share param) $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Speed(tokens/s/gpu) 13400 12500 10600 13800 14600 13000 15000 10500
GPU Memory(GB) 65 65 69 69 61 75 66 75
Table 2: Comparison of different training configurations.
5 Pretraining
The pretraining corpus employed in our study predominantly consists of Chinese texts and is entirely derived from open-source datasets. It includes prominent datasets such as SkyPile-150B  (Wei et al., 2023), Wanjuan1.0 (He et al., 2023), Wikipedia-cn, as well as diverse chat data from multiple sources and Starcode. Further details of these datasets are provided in Appendix A.
Figure 2: Pretraining Data Distribution
To ensure consistency across the data, we initially standardized all datasets into a uniform format. Subsequently, we utilized Alibaba’s open-source tool, Data-Juicer, to meticulously filter and transcribe both the text and code data. We utilized a total of 21 text processing operators, as described in Appendix C, and 13 code processing operators, as listed in Appendix D. The final step involved employing the tokenizer from the Qwen1.5. This tokenizer was instrumental in converting the entire corpus into token-ids, which were then merged and segmented into manageable chunks. The distribution of the pretraining data across different domains is illustrated in Figure 2. Despite these efforts, our data preprocessing workflow still exhibits certain limitations, particularly the lack of balance in the proportion of texts from different domains within the pre-training corpus.
Over a span of 30 days, Steel-LLM was trained through 1.07 million steps. The initial 200,000 steps utilized NVIDIA 8 A100-80G GPUs for training, while the remaining steps employed 8 H800-80G GPUs. The loss curve is shown in Appendix E. We set the maximum sequence length to 2048, the batch size to 8, and the number of gradient accumulation steps to 8. Consequently, Steel-LLM was trained on approximately one trillion tokens. We employ the AdamW optimizer (Kingma & Ba (2017)) with hyper-parameters set to $\beta_{1}$= 0.9, $\beta_{2}$= 0.95, and weight_decay = 0.05. The maximum learning rate is set to $3\times 10^{-4}$, and the gradient clipping norm is set to 1.0. We employed a cosine-annealing learning rate schedule with 2,000 warmup steps, such that the final learning rate is 0.
6 Post training
6.1 Supervised Finetuning
There has been substantial discussion and some debate among researchers regarding the volume of datasets required for supervised fine-tuning. For instance, Zhou et al. (2023) advocates for selecting a few hundred to a few thousand high-quality data points for fine-tuning, while Ji et al. (2023) suggests that using a significantly larger amount of data can yield better results. Given the capabilities of our base model, we determined that augmenting the fine-tuning phase with additional data was necessary. This strategy not only aids the model in mastering conversational techniques but also enriches it with supplementary knowledge. Consequently, We curated the following fine-tuning datasets: Infinity-Instruct Chinese data (BAAI, 2024) (approximately 700,000 entries, with all English data removed), multiple-choice questions from the Wanjuan-cn dataset (He et al., 2023) which tested in both CoT and non-CoT formats, Ruozhiba data (Ruozhiba, 2024), self-awareness data with template modified from (Team, 2024a), and three English datasets: Code-Feedback (Zheng et al., 2025), WebInstructSub (Yue et al., 2024) and OpenHermes-2.5 (Teknium, 2023). Further details are listed in Appendix B.
We fine-tuned the pre-trained Steel-LLM for approximately 4 epochs using the SFT dataset. The global batch size was set to 256, the maximum learning rate was set to $2\times 10^{-5}$, and a cosine-annealing learning rate schedule was employed. The loss curve is shown in Appendix F. Below, we present a series of ablation studies and evaluations to analyze the impact of different fine-tuning strategies on model performance.
6.2 Ablation Studies and Evaluation
To systematically evaluate the effectiveness of our fine-tuning approach, we conducted several experiments with varying data compositions and formats. The results are summarized in Table 3, and the key findings are discussed below.
Experiment CEVAL Acc. CMMLU Acc. MMLU Acc.
Full Infinity-Instruct + Wanjuan MCQ 32.35 26.32 25.50
700K Chinese Infinity-Instruct + Wanjuan MCQ 38.57 33.48 23.26
Chinese + 100% English Data 39.21 33.20 26.73
Chinese + 20% English Data (Balanced) 40.43 35.86 26.75
Chinese + 20% English + English MCQ 41.90 36.08 30.82
Table 3: Performance of different fine-tuning strategies.
Experiment 1: Full Infinity-Instruct + Wanjuan MCQ
In this experiment, we fine-tuned the model using the entire Infinity-Instruct dataset (approximately 7,000,000 entries) and the full Wanjuan multiple-choice dataset. While the model’s performance on CEVAL improved with increasing fine-tuning steps, the accuracy plateaued at around 32% after 18,000 steps. This suggests that even though the Wanjuan data was included in the pretraining phase, the model still benefited from additional fine-tuning, likely due to its relatively small size.
Experiment 2: 700K Chinese Infinity-Instruct + Wanjuan MCQ
To address the issue of language mismatch, we filtered out the English data from Infinity-Instruct and retained only the 700,000 Chinese entries. This adjustment significantly improved the model’s performance, achieving 38% accuracy on CEVAL and 33% on CMMLU. This result highlights the importance of aligning the fine-tuning data distribution with the pretraining phase.
Experiment 3: Chinese + 100% English Data
To explore the impact of English data on the model’s multilingual capabilities, we fine-tuned the model using a balanced mix of Chinese and English data (340,000 entries each). While the model’s performance on Chinese benchmarks remained stable, its performance on MMLU improved slightly, indicating that English data can complement the model’s existing knowledge without degrading its Chinese capabilities.
Experiment 4: Chinese + 20% English Data (Balanced)
In this experiment, we fine-tuned the model with a data distribution that closely matched the pretraining phase (80% Chinese and 20% English). This approach not only improved the model’s performance on Chinese benchmarks (CEVAL: 39.21% → 40.43%; CMMLU: 33.2% → 35.86%) but also maintained its performance on MMLU. This suggests that maintaining a balanced data distribution during fine-tuning is crucial for preserving the model’s multilingual capabilities.
Experiment 5: Chinese + 20% English + English MCQ
To further enhance the model’s reasoning abilities, we introduced additional English multiple-choice questions from datasets such as OpenBookQA, AI2 ARC, and LogiQA. This experiment resulted in a noticeable improvement in MMLU accuracy (26.75% → 30.82%), while the performance on Chinese benchmarks remained stable. This indicates that incorporating domain-specific question-answering data can enhance the model’s reasoning capabilities without compromising its existing knowledge.
6.3 Learning from Human Preferences
To align Steel-LLM with human preferences, we employ the Direct Preference Optimization (DPO)Rafailov et al. (2024) algorithm and sorted response pairs for model optimization.
In the preference dataset, the ratio of Chinese to English stands at 4:1, consistent with that in the pre-training phase. The Chinese dataset is derived from ultrafeedback-chinese [^1] https://huggingface.co/datasets/opencsg/UltraFeedback-chinese/, and the English dataset is derived from ultrafeedback-binarized-preferences [^2] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences. Both the reference model and the objective model of the DPO algorithm are initialized with the SFT version of Steel-LLM. We conducted training on the data for 3 epochs. The global batch size was set to 128, the maximum learning rate was set to $5\times 10^{-6}$, the pref_beta set to 0.1, and a cosine-annealing learning rate schedule was employed. The loss curve is shown in Appendix G.
6.4 Discussion
Our ablation studies reveal several key insights:
  • Data Distribution Matters: Fine-tuning with a data distribution that closely matches the pretraining phase leads to better performance on both Chinese and English benchmarks.
  • Small Models Benefit from Additional Data: Even for small models, fine-tuning with a larger dataset can improve performance, particularly when the pretraining data is limited.
  • Balanced Multilingual Fine-Tuning: Incorporating a small proportion of English data during fine-tuning can enhance the model’s multilingual capabilities without degrading its performance on Chinese tasks.
  • Exam-Style Data Enhances Performance: Including domain-specific question-answering data, such as multiple-choice questions, can improve the model’s reasoning abilities and overall benchmark performance.
To contextualize the performance of Steel-LLM, we compare it with several state-of-the-art models on the CEVAL and CMMLU benchmarks, as shown in Table 4. Our best-performing model, Steel-LLM-Chat, achieves competitive results, outperforming models of similar scale such as Tiny-Llama-1.1B and Gemma-2b-it, and approaching the performance of larger models like CT-LLM-SFT-2B. While Steel-LLM-Chat does not yet match the performance of significantly larger models like Qwen1.5-1.8B-Chat or Qwen-7B, its results demonstrate the effectiveness of our resource-efficient approach, particularly given the limited computational resources used for training.
Model CEVAL CMMLU
Tiny-Llama-1.1B (Zhang et al., 2024c) 25.02 24.03
MiniCPM-1.2B (min, 2024) 49.14 46.81
Qwen1.5-1.8B-Chat (Bai et al., 2023b) 56.84 54.11
Phi2(2B)(Abdin et al., 2023) 23.37 24.18
Gemma-2b-it (Gemma Team et al., 2024) 32.30 33.07
CT-LLM-SFT-2B (Du et al., 2024) 41.54 41.48
ChatGLM-6B (GLM et al., 2024) 38.90 37.48
Llama2-7B Touvron et al. (2023b) 32.42 31.11
OLMo-7B (Groeneveld et al., 2024b) 35.18 35.55
Gemma-7B (Gemma Team et al., 2024) 42.57 44.20
MAP-Neo-7B (Zhang et al., 2024a) 56.97 55.01
Llama2-13B Touvron et al. (2023b) 37.32 37.06
Steel-LLM-1B-Chat 41.90 36.08
Steel-LLM-1B-Chat-DPO 42.04 36.04
Table 4: Performance comparison of models on CEVAL and CMMLU benchmarks.
In conclusion, our fine-tuning approach demonstrates that careful dataset selection and composition can significantly enhance the performance of small language models like Steel-LLM. By balancing the inclusion of diverse data types and maintaining alignment with the pretraining distribution, we achieved competitive results on both Chinese and English benchmarks, positioning Steel-LLM as a strong contender among resource-efficient LLMs.
7 Conclusion
This paper introduces Steel-LLM, a fully open-source Chinese-centric language model developed with limited computational resources, achieving competitive performance on benchmarks such as CEVAL (41.90%) and CMMLU (36.08%). By leveraging innovative techniques like Soft Mixture of Experts and enhanced feed-forward networks, along with systematic training optimizations, we demonstrate that high-quality LLMs can be built efficiently. Our work provides complete transparency, releasing the training pipeline, datasets, and intermediate checkpoints, offering practical guidance for small-scale LLM development. All resources are made publicly available to foster collaboration and advance accessible language technologies.
Appendix A Pretraining Data Detailed Description
Dataset Description
SkyPile-150B (Wei et al., 2023) Consisting of approximately 150 billion tokens and 620 gigabytes of cleaned text data from 233 million web pages, with rigorous filtering and deduplication to ensure quality and mitigate sensitive and biased information.
Wanjuan1.0 (He et al., 2023) Composed of processed data from various sources, including web pages, encyclopedias, books, patents, textbooks, and exam questions, with a total volume of data exceeding 500 million documents, amounting to over 1TB (roughly split equally between Chinese and English data) and has undergone meticulous cleaning, deduplication, and value alignment.
Wikipedia-cn Based on the July 20th, 2023 Chinese Wikipedia dump, retains 254,574 high-quality entries after filtering out special types, low-quality, sensitive, and controversial content, and includes conversions between simplified and traditional Chinese.
Baidu Baike Consisting of 5,630,000 uncleaned entries from Baidu Baike, with a total size of approximately 17GB.
Baidu QA Including 1.5 million high-quality encyclopedia questions and answers, spanning 492 categories, with 434 categories occurring at least 10 times, suitable for training intelligent Q&A systems
Zhihu QA Including 1 million entries, with 1.5GB in size.
BELLE (BELLEGroup, 2023) Train_2M_CN and train_3.5M_CN, generated by ChatGPT, contain 2 million and 3.5 million dialogue entries respectively, and were both used in this project. Note that these datasets are unverified and may contain errors.
Moss (Sun et al., 2024) Containing 1.1 million Chinese and English multi-turn dialogue entries.
Firefly (Yang, 2023) Comprising 1.15 million entries covering 23 common Chinese NLP tasks and includes culturally relevant data such as couplets, poetry, classical Chinese translations, prose, and Jin Yong novels, resulting in a total of 1.15 million entries.
Starcode (Li et al., 2023) Including 783GB of code across 86 programming languages, with 54GB of GitHub Issues, 13GB of Jupyter notebooks, and 32GB of GitHub commits. Our project used only the C++, Python, and Java data.
Table 5: Pretraining Data Detailed Description
Appendix B Supervised Finetuning Data Detailed Description
Dataset Description
Infinity-Instruct-7M (BAAI, 2024) A large-scale, high-quality instruction dataset with only 0.7M of Chinese data used in this project.
Wanjuan1.0 (He et al., 2023) Consistent with the one used during the pre-training stage, but with the Chinese choice question data repurposed for fine-tuning.
Ruozhiba (Ruozhiba, 2024) Questions from Baidu Tieba “Ruozhiba” were answered by GPT-4, then manually reviewed and edited for formatting errors and improved responses.
Self-awareness Dataset (Team, 2024a) Consisting of various “Who are you?” questions from the EmoLLM project templates.
Code-Feedback (Zheng et al., 2025) A code SFT dataset consists of 66,000 entries from various open-source code datasets and LeetCode, after undergoing a series of filtering and selection processes.
WebInstructSub (Yue et al., 2024) Containing 2.33 million SFT entries across fields such as mathematics, physics, biology, chemistry, and computer science.
OpenHermes-2.5 (Teknium, 2023) Consisting of samples synthesized by large models and chat samples, filtered from open-source data like Airoboros, ChatBot Arena, and Evol Instruct, totaling 1 million entries.
Table 6: Supervised Finetuning Data Detailed Description
Appendix C Data Juicer Operators Used for Text Processing
Operator Description Note
chinese_convert_mapper Converts Chinese between Traditional Chinese, Simplified Chinese and Japanese Kanji Mode: t2s (tradition to simple)
clean_email_mapper Removes email information -
clean_html_mapper Removes HTML tags and returns plain text of all the nodes -
clean_ip_mapper Removes IP addresses -
clean_links_mapper Removes links, such as those starting with http or ftp -
clean_copyright_mapper Removes copyright notice at the beginning of code files (must contain the word copyright) -
expand_macro_mapper Expands macros usually defined at the top of TeX documents -
fix_unicode_mapper Fixes broken Unicodes -
punctuation_normalization_mapper Normalizes various Unicode punctuations to their ASCII equivalents -
remove_repeat_sentences_mapper Remove repeat sentences in text samples Ignore special char- acter and sentences shorter than 2 will not be deduplicated
remove_specific_chars_mapper Removes any user-specified characters or substrings -
whitespace_normalization_mapper Normalizes various Unicode whitespaces to the normal ASCII space (U+0020) -
alphanumeric_filter Keeps samples with alphanumeric ratio within the specified range [0.0, 0.9]
average_line_length_filter Keeps samples with average line length within the specified range [10, 150]
character_repetition_filter Keeps samples with char-level n-gram repetition ratio within the specified range [0.0, 0.4]
maximum_line_length_filter Keeps samples with maximum line length within the specified range 1000
perplexity_filter Keeps samples with perplexity score below the specified threshold 1500
special_characters_filter Keeps samples with special-char ratio within the specified range [0.0, 0.25]
text_length_filter Keeps samples with total text length within the specified range [10, 100000]
word_repetition_filter Keeps samples with word-level n-gram repetition ratio within the specified range [0.0, 0.5]
document_simhash_deduplicator Deduplicates samples at document-level using SimHash Tokenization:space; window_size:6; num_blocks:6; hamming_distance:4; lowercase:true
Table 7: Data Juicer Operators Used for Text Processing
Appendix D Data Juicer Operators Used for Code Processing
Operator Description Note
clean_copyright_mapper Removes copyright notice at the beginning of code files (must contain the word copyright) -
clean_email_mapper Removes email information -
clean_links_mapper Removes links, such as those starting with http or ftp -
fix_unicode_mapper Fixes broken Unicodes -
punctuation_normalization_mapper Normalizes various Unicode punctuations to their ASCII equivalents -
alphanumeric_filter Keeps samples with alphanumeric ratio within the specified range [0.546, 3.65]
average_line_length_filter Keeps samples with average line length within the specified range [10, 150]
character_repetition_filter Keeps samples with char-level n-gram repetition ratio within the specified range 0.36
maximum_line_length_filter Keeps samples with maximum line length within the specified range 1000
text_length_filter Keeps samples with total text length within the specified range 96714
words_num_filter Keeps samples with word count within the specified range [20,6640]
word_repetition_filter Keeps samples with word-level n-gram repetition ratio within the specified range [10, 0.357]
document_simhash_deduplicator Deduplicates samples at document-level using SimHash Tokenization:space; window_size:6; num_blocks:6; hamming_distance:4; lowercase:true
Table 8: Data Juicer Operators Used for Code Processing
Appendix E Pre-training Loss Curve
The loss curve during the pre-training stage is shown in Figure 3. The initial 200,000 steps utilized NVIDIA 8 A100-80G GPUs for training, while the remaining steps employed 8 H800-80G GPUs.
Figure 3: Pre-training loss curve for Steel-LLM
Appendix F Supervised Finetuning Loss Curve
Figure 4: Supervised Fine-tuning loss curve for Steel-LLM
Appendix G Direct Preference Optimization Loss Curve
Figure 5: Direct Preference Optimization loss curve for Steel-LLM

### v2

@@
 Several initiatives, such as LLM360’s Amber (Liu et al., 2023; Tan et al., 2024), M-A-P’s Neo (Zhang et al., 2024b), and AI2’s OLMO series (Groeneveld et al., 2024a; OLMo et al., 2024; Muennighoff et al., 2024), have addressed these limitations by releasing complete training pipelines, datasets, and intermediate checkpoints.
-While these contributions are invaluable, their development typically demands extensive computational resources, making them inaccessible to smaller research teams or individual practitioners.
+While these contributions are invaluable, their development typically requires extensive computational resources, making them inaccessible to smaller research teams or individual practitioners.
 This creates a significant gap for fully transparent and resource-efficient LLMs tailored for smaller-scale research efforts, particularly in non-English languages such as Chinese.
@@
 The model architecture incorporates innovative features such as Soft Mixture of Experts (Soft MoE) and an enhanced Feed-Forward Network, optimizing performance within resource constraints.
-Our training framework, modified from TinyLlama (Zhang et al., 2024c), includes several optimizations for efficiency and usability, such as improved model loading, training progress restoration, and the ability to append data during the training process.
+Our training framework, modified from TinyLlama (Zhang et al., 2024c), includes several optimizations for efficiency and usability, such as improved model loading, restoration of training progress, and the ability to append data during the training process.
 The contributions of this work are as follows:
@@
   • Complete Transparency: We offer complete transparency in our development process, including the release of our training pipeline, dataset, model architecture, and intermediate checkpoints.
-This facilitates reproducibility and enables further research by the broader community.
+This facilitates reproducibility and allows for further research by the broader community.
   • Practical Guidance for Small-Scale Research: We provide detailed insights into our model architecture, training framework, and data preparation process, offering practical guidance for researchers and practitioners with limited resources.
@@
 2 Related Works
-Recent developments in open-source LLMs have varied widely in terms of transparency and accessibility.
+Recent developments in open source LLM have varied widely in terms of transparency and accessibility.
 Many models, such as Qwen (Bai et al., 2023a; Yang et al., 2024), Llama (Touvron et al., 2023b), Deepseek (DeepSeek-AI et al., 2024a), Gemma  (Team et al., 2024), InternLM  (Cai et al., 2024), Mixtral  (Jiang et al., 2023), Yi  (AI et al., 2024), GLM  (GLM et al., 2024), have been released with only the final model checkpoints and weights, while withholding crucial details such as training data, codes, and intermediate checkpoints.
@@
 Under the condition of the same parameter scale, the sparse MoE model activates fewer parameters than a dense model, thus reducing the number of FLOPs.
-Steel-LLM is trained using only on 8 GPUs with limited GPU memory.
+Steel-LLM is trained using only 8 GPUs with limited GPU memory.
 For a model featuring sparse MoE structure, all parameters must be loaded into the GPU memory.
@@
 Soft MoE is fully differentiable.
-Therefore, there is no need to consider problems such as expert unbalance that exist in sparse MoE.
+Therefore, there is no need to consider problems such as expert imbalance that exist in sparse MoE.
 We denote the input tokens for a single sequence as $\mathbf{X}\in\mathbb{R}^{m\times d}$, where $m$ represents the number of tokens and $d$ denotes their dimensions.
@@
 3.2 Enhanced Feed-Forward Network
-Since the Steel-LLM employs the soft MoE approach, within this framework, the Feed-Forward Network (FFN) effectively represents an expert.
+Since Steel-LLM employs the soft MoE approach, within this framework, the Feed-Forward Network (FFN) effectively represents an expert.
 In a vanilla Transformer, The FFN comprises two layers of Multi-Layer Perceptrons (MLPs).
@@
 We normalize the inputs of self-attention layers and FFN layers with RMSNorm(Zhang & Sennrich (2019)).
-The hyperparameters of Steel-LLM are presented in Table 1. We employed the Qwen1.5 tokenizer (Team (2024b)), which utilizes byte pair encoding (BPE).
+The hyperparameters of Steel-LLM are presented in Table 1. We employ the Qwen1.5 tokenizer (Team (2024b)), which utilizes byte pair encoding (BPE).
 4 Training Framework
@@
 Additionally, to avert the inadvertent repeated addition of data files to the training process, Steel-LLM has incorporated a function that utilizes a hash value, specifically the MD5(Rivest (1992)) hash algorithm, to detect data content duplication.
-To imporve training speed and conserve GPU memory, during the training process, we utilized techniques such as bfloat16 mixed-precision(Kalamkar et al.
+To improve training speed and conserve GPU memory, during the training process, we utilized techniques such as bfloat16 mixed-precision(Kalamkar et al.
 (2019)), Fully Sharded Data Parallel (FSDP, Zhao et al.
@@
 We utilized a total of 21 text processing operators, as described in Appendix C, and 13 code processing operators, as listed in Appendix D.
-The final step involved employing the tokenizer from the Qwen1.5.
+The final step involved employing the tokenizer from Qwen1.5.
 This tokenizer was instrumental in converting the entire corpus into token-ids, which were then merged and segmented into manageable chunks.
@@
 Baidu QA Including 1.5 million high-quality encyclopedia questions and answers, spanning 492 categories, with 434 categories occurring at least 10 times, suitable for training intelligent Q&A systems
-Zhihu QA Including 1 million entries, with 1.5GB in size.
-BELLE (BELLEGroup, 2023) Train_2M_CN and train_3.5M_CN, generated by ChatGPT, contain 2 million and 3.5 million dialogue entries respectively, and were both used in this project.
+Zhihu QA Including 1 million entries of questions and answers, with 1.5GB in size.
+BELLE (BELLEGroup, 2023) Including train_2M_CN and train_3.5M_CN, which are generated by ChatGPT, containing 2 million and 3.5 million dialogue entries respectively, and both used in this project.
 Note that these datasets are unverified and may contain errors.
 Moss (Sun et al., 2024) Containing 1.1 million Chinese and English multi-turn dialogue entries.
-Firefly (Yang, 2023) Comprising 1.15 million entries covering 23 common Chinese NLP tasks and includes culturally relevant data such as couplets, poetry, classical Chinese translations, prose, and Jin Yong novels, resulting in a total of 1.15 million entries.
+Firefly (Yang, 2023) Comprising 1.15 million entries which cover 23 common Chinese NLP tasks and include culturally relevant data such as couplets, poetry, classical Chinese translations, prose, and Jin Yong’s novels, resulting in a total of 1.15 million entries.
 Starcode (Li et al., 2023) Including 783GB of code across 86 programming languages, with 54GB of GitHub Issues, 13GB of Jupyter notebooks, and 32GB of GitHub commits.
@@
 WebInstructSub (Yue et al., 2024) Containing 2.33 million SFT entries across fields such as mathematics, physics, biology, chemistry, and computer science.
-OpenHermes-2.5 (Teknium, 2023) Consisting of samples synthesized by large models and chat samples, filtered from open-source data like Airoboros, ChatBot Arena, and Evol Instruct, totaling 1 million entries.
+OpenHermes-2.5 (Teknium, 2023) Consisting of samples synthesized by large models and chat data, filtered from open-source data like Airoboros, ChatBot Arena, and Evol Instruct, totaling 1 million entries.
 Table 6: Supervised Finetuning Data Detailed Description